OpenStudy (anonymous):

differentiate log x 3^x

OpenStudy (anonymous):

is that \[3^xlog x\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

you will need to use the product rule for this. y'=u'v+v'u are you familiar with this?

OpenStudy (anonymous):

how would you dash 3^x?

OpenStudy (anonymous):

here is the method, a little trick is used to make it easier http://www.math.com/tables/derivatives/more/b%5Ex.htm

OpenStudy (anonymous):

i don't understand =(

OpenStudy (anonymous):

okay lets just look as the derivative of 3^x \[3^x \frac{d}{dx} = e^{ln3^x} \frac{d}{dx}\] \[=e^{xln3} \frac{d}{dx}\] Now let u =xln3 then x=u/ln3 \[=e^{u} ln3\frac{d}{du}\]\[=ln3e^{u} \]\[=ln3e^{xln3} \]\[=ln3e^{ln3^x} \]\[=(ln3)3^x \]

OpenStudy (anonymous):

Differential of a^x= a^x lna